OpenStudy (anonymous):

2(6d+3)=18-3(16-3d) what is d

OpenStudy (akashdeepdeb):

Use the Distributive property! Like this... 12d + 6 = 18 - 48 + 9d Now solve the simultaneous equation! :)

OpenStudy (anonymous):

Akashi I did can't solve it

OpenStudy (austinl):

You need to combine like terms.

OpenStudy (akashdeepdeb):

Just take the variables to one side and the constants to the other like this.. 12d + 6 = 18 - 48 + 9d 12d - 9d = -30 - 6 3d = -36 Got it? :)